# Gaussian Scenes: Pose-Free Sparse-View Scene Reconstruction using Depth-Enhanced Diffusion Priors

### v2

@@ Appendix D Qualitative Comparison with ReconFusion / CAT3D
Figure 13: Qualitative comparison of GScenes with ReconFusion and CAT3D (posed techniques). Despite being a pose-free pipeline built with weaker diffusion priors, our method achieves competitive NVS quality with SOTA sparse-view reconstruction techniques. No image available for CAT3D in the last row, hence kept blank.
@@ Appendix E Additional Qualitative Results
[…]
@@
[…]
@@ Appendix F Ablation Studies
Appendix G Limitations & Future Work
Table 4: Pose estimation accuracy with GScenes. SfM-poses for training views estimated from the full observation set are used as ground truth. We report errors in camera rotation and translation using Absolute Trajectory Error (ATE) and Relative Pose Error (RPE) as in  Bian et al. (2023). Despite the DUSt3r initialization of camera extrinsics and subsequent pose optimization, there are large errors w.r.t COLMAP poses of the dense observation set, harming test-pose alignment and subsequent NVS quality.

### v3

@@ Appendix D Qualitative Comparison with ReconFusion / CAT3D
Figure 13: Qualitative comparison of GScenes with ReconFusion and CAT3D (posed techniques). Despite being a pose-free pipeline built with weaker diffusion priors, our method achieves better (flowers, treehill) or similar (plant) NVS quality compared to ReconFusion on three out of 4 examples above. We are slightly worse compared to closed-source CAT3D (although their results cannot be validated), which uses a closed-source, stronger video diffusion prior. No image available for CAT3D in the last row, hence kept blank.
@@ Appendix E Additional Qualitative Results
MASt3R + 3DGS RegNeRF ´ DiffusioNeRF ´ ZeroNVS* ´ COGS Ours Ground Truth
@@
MASt3R + 3DGS ZeroNVS* ´ DiffusioNeRF ´ COGS Ours Ground Truth
@@ Appendix F Ablation Studies
Appendix G Evaluating DiffusioNeRF in a pose-free setting
Since the performance of ReconFusion and CAT3D in a pose-free setting cannot be ascertained due to a lack of open-source code or models, we instead evaluate DiffusioNeRF, an open-source posed reconstruction technique, on 3 scenes from the DL3DV-benchmark in the same pose-free evaluation protocol as COGS, MASt3R+3DGS, and GScenes. We use noisy poses estimated by MASt3R from sparse inputs as training poses instead of ground truth COLMAP poses (posed setting) and show the difference in NVS quality on test views in Fig 16. From this experiment, one can conclude that NVS quality achieved with posed techniques does not directly translate to similar performance in a pose-free setting.
DiffusioNeRF (posed) DiffusioNeRF (pose-free) Ground Truth
3 views Refer to caption Refer to caption Refer to caption
6 views Refer to caption Refer to caption Refer to caption
3 views Refer to caption Refer to caption Refer to caption
Figure 16: DiffusioNeRF in a pose-free setting. We evaluate DiffusioNeRF using poses estimated by MASt3R instead of ground truth colmap poses estimated from dense views. The drop in NVS quality demonstrates that a posed reconstruction technique cannot be trivially adapted to the pose-free setting.
Appendix H Effect of Generative Priors on Training Views
Table 4: Effect of Generative Priors on Training Views. We pick the Square scene from DL3DV-10K and its 3, 6, 9 view splits to compare reconstruction performance of GScenes with our baseline MASt3R + 3DGS on the training views. Noise and 3D inconsistencies introduced from the generated novel views cause a minor drop-off in the reconstruction quality of training views.
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
3-view 6-view 9-view 3-view 6-view 9-view 3-view 6-view 9-view
MASt3R + 3DGS 23.34 22.32 22.96 0.73 0.62 0.63 0.20 0.21 0.23
GScenes 22.44 21.52 21.70 0.75 0.57 0.56 0.18 0.25 0.26
MASt3R + 3DGS (3) GScenes (3) MASt3R + 3DGS (6) GScenes (6) MASt3R + 3DGS (9) GScenes (9) Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 17: Effect of Generative Priors on Training Views.
In this section, we analyze the effect of using generative priors for novel view synthesis on the reconstruction quality of the sparse training views of a 3D scene. To ensure 3D consistency and limited artifacts in the reconstructed scene, our diffusion model integrates camera, context, and confidence map conditioning, and the autoregressive 2D-3D distillation process (Sec  2.6) progressively resolves inconsistencies across different novel views. The differences that cannot be reconciled result in artifacts like floaters, which also negatively impact the reconstruction quality of the input views. In Tab 4 and Fig 17, we pick the Square scene from DL3DV-10K benchmark and its 3, 6, and 9 view splits to compare the performance of GScenes with our baseline reconstruction engine - MASt3R + 3DGS on reconstruction of training views. As we can observe, while there is minor degradation both qualitatively and quantitatively due to distillation of generative priors, the original 3DGS objective for training views during autoregressive scene reconstruction ensures that the reconstructed sparse input views still faithfully represent the 3D scene.
Appendix I Limitations & Future Work
Table 5: Pose estimation accuracy with GScenes. SfM-poses for training views estimated from the full observation set are used as ground truth. We report errors in camera rotation and translation using Absolute Trajectory Error (ATE) and Relative Pose Error (RPE) as in  Bian et al. (2023). Despite the MASt3r initialization of camera extrinsics and subsequent pose optimization, there are large errors w.r.t COLMAP poses of the dense observation set, harming test-pose alignment and subsequent NVS quality.